OpenStudy (anonymous):

A fair octahedral die is rolled - what is the probability that after five rolls at least one number has been obtained twice?

OpenStudy (anonymous):

ok lets call A = after five rolls at least one number has been obtained twice and lets call A' = after five rolls no numbers were repeated P(A) = 1 - P(A') do you agree?

OpenStudy (anonymous):

Got that

OpenStudy (anonymous):

ok so lets find P(A') as it might be easier what is the probability i have no repeated numbers after 1 roll? well its 1 as i cant have repeats on only one roll what is the probability that i have no repeats after two rolls? well whichever number we got on the first roll we cant have again, so this time round we have \[\frac{7}{8}\] remember this relies on the first case being true: \[1 \times \frac{7}{8}\] now we have rolled the die twice and have used up two numbers, we have 6 left that are allowed on the third roll: \[1 \times \frac{7}{8} \times \frac{6}{8}\] i'm going to change the 1 to 8/8 because it looks nicer and you can see the pattern emerge: \[\frac{8}{8} \times \frac{7}{8} \times \frac{6}{8}\] so thats after three rolls, can you tell me what P(A') is after five rolls?

OpenStudy (anonymous):

8/8 * 7/8 * 6/8 * 1/8 * 1/8 ?

OpenStudy (anonymous):

hmm not quite, i'll use a specific example lets say i roll the die and get a 1 , then on the next go i am allowed any of the following: {2,3,4,5,6,7,8} so the probability of no repeats on the second roll is 7/8 times the probability of no repeats on the first go lets say i roll and get a 2, now we are allowed any of the following: {3,4,5,6,7,8} so the probability for no repeats on this go is 6/8 times the probability of no repeats by the second go: 6/8 x 7/8 x 1 so now i roll and get a 3, what numbers are left? well: {4,5,6,7,8} so 5/8 x 6/8 x 7/8 x 1 what numbers am i allowed if i now roll a 4 and dont want a repeated value?

OpenStudy (anonymous):

5, 6, 7, 8

OpenStudy (anonymous):

yep, so now we're on the fifth roll and the probability of no repeats so far is 4/8 x 5/8 x 6/8 x 7/8 x 8/8

OpenStudy (anonymous):

so calculate this value, its P(A') and then P(A) = 1- P(A')

OpenStudy (anonymous):

There are 8^5 outcomes when we throw a fair die having 8 sides. There are \[\frac {8!} {3! 5!} \] ways to choose five distinct numbers from 1 to 8. Since order is not important. There are \[5!\frac {8!} {3! 5!}=\frac {8!} {3! } \] The probability of no repeats is \[ \frac {\frac {8!} {3! }} {8^5}=\frac{105}{512} \] Which is exactly what you get by what @eigenschmeigen was trying to guide you through.